OpenStudy (anonymous):

solve for "y" 8-2y = 9y+(y-2)

OpenStudy (anonymous):

first open the brackets then bring all the y to one side

OpenStudy (anonymous):

@nphuongsun93 can you explain?

OpenStudy (anonymous):

8 - 2y = 9y + (y - 2) open brackets 8 - 2y = 9y + y - 2 8 - 2y = 10y - 2 + 2y and + 2 both sides 10 = 12y

OpenStudy (anonymous):

@nphuongsun93 oops forgot a number! 8-2y=9y+ 3 (y-2) got it thanks!

OpenStudy (anonymous):

oh just times 3 into (y-2) when you open the brackets

OpenStudy (anonymous):

THANK YOU!